# A Survey on Self-Supervised Learning for Non-Sequential Tabular Data

### v1

@@ Abstract
Self-supervised learning (SSL) has been incorporated into many state-of-the-art models in various domains, where SSL defines pretext tasks based on unlabeled datasets to learn contextualized and robust representations. Recently, SSL has been a new trend in exploring the representation learning capability in the realm of tabular data, which is more challenging due to not having explicit relations for learning descriptive representations. This survey aims to systematically review and summarize the recent progress and challenges of SSL for non-sequential tabular data (SSL4NS-TD). We first present a formal definition of NS-TD and clarify its correlation to related studies. Then, these approaches are categorized into three groups – predictive learning, contrastive learning, and hybrid learning, with their motivations and strengths of representative methods within each direction. On top of this, application issues of SSL4NS-TD are presented, including automatic data engineering, cross-table transferability, and domain knowledge integration. In addition, we elaborate on existing benchmarks and datasets for NS-TD applications to discuss the performance of existing tabular models. Finally, we discuss the challenges of SSL4NS-TD and provide potential directions for future research. We expect our work to be useful in terms of encouraging more research on lowering the barrier to entry SSL for the tabular domain and improving the foundations for implicit tabular data.
@@ 1 Introduction
Supervised learning has shown outstanding performance on various machine learning tasks; however, its main hurdles lie in heavily depending on expensive human annotations and generalization bottlenecks. With the scaling of the accessibility to unlabeled data, self-supervised learning (SSL) has witnessed its generic and robust ability; that is, learning contextualized information from correlations within the data. A line of studies has shown that SSL is able to push new boundaries in various domains, including text Hoffmann et al. (2022); Li et al. (2023b), vision Li et al. (2023a); Woo et al. (2023), and speech Baevski et al. (2021); Wu (2023). In addition, SSL illustrates strong generalizability, enabling models to adapt to tasks with limited labeled records and even unseen tasks Balestriero et al. (2023). The key advantage of SSL is that it reduces efforts in annotating a large amount of data, while further providing generalization ability.
[…]
@@
Previous SSL methods highly relied on the unique structure of the domain datasets, such as spatial relationships in images, semantic relationships in text, and vocal relationships in speech. In contrast, tabular data have no explicit relation between each feature, and may be completely different across various tabular datasets. Figure 1 describes the overall pipeline of learning representations from tabular data. To opt for the learning strategy, in contrast to only using supervised learning requiring task-specific labels during the training stage, SSL is further leveraged to learn task-agnostic representations from pretext tasks, creating labels explicitly (e.g., predictive learning) and/or implicitly (e.g., contrastive learning). The model is expected to learn universal representations from unlabeled tabular datasets and accordingly adapt effectively to different downstream tasks including both classification and regression problems. Generally, existing techniques with SSL for tabular data can be grouped into sequential and non-sequential tabular types for representation learning.
@@
This survey focuses on SSL for non-sequential tabular data (SSL4NS-TD) for two reasons. First, SSL for sequential data commonly used in fields such as recommendation systems has been widely adopted by recurrent- and Transformer-based techniques based on the temporal-ordered composition, which is already discussed in the recent survey papers Badaro et al. (2023); He et al. (2023). On the flip side, non-sequential tabular data (NS-TD) make it more challenging to define explicit structures with pretext tasks, which has led to tremendous efforts in recent years but has not yet been discussed. Accordingly, it inspires us to provide a systematic review of recent SSL4NS-TD approaches to discuss their motivations and self-supervised objectives. Second, whether deep learning models are superior to machine learning models in the NS-TD problems remains an active discussion Grinsztajn et al. (2022); McElfresh et al. (2023). However, SSL showcases its robustness in not only full-labeled data but also only a few labeled records, where machine learning models fall short in such low-resource scenarios. It is thus urgent to summarize the efforts of learning contexts with self-supervision for NS-TD to sensitize the community to the current progress and open with more discussion.
This survey paper will be a contribution to both researchers and industrial practitioners with NS-TD applications, e.g., finance and healthcare. The SSL-enabled solutions for these applications have a common ground in learning contextualized representations for tabular data with multifaceted learning strategies. Moreover, the comprehensive discussions of existing NS-TD benchmarks are able to stimulate a more thorough empirical evaluation of contributions. This survey presents an up-to-date paper survey with a paper list that will be continuously updated [^1] A detailed list is at https://github.com/wwweiwei/awesome-self-supervised-learning-for-tabular-data., and provides an in-depth discussion on related studies of SSL4NS-TD.
@@
The remainder of this review is structured as follows. Section 2 first introduces the problem of NS-TD and proposes a novel taxonomy of SSL4NS-TD consisting of three learning strategies as summarized in Table 1. The detailed discussion of the achievements, the downstream tasks, and code links (if available) of the three directions are summarized in Sections 3, 4, and 5 respectively. Section 6 briefly introduces application issues (i.e., automatic data engineering, cross-table transferability, and domain knowledge integration) in practice and recent related work. Subsequently, Section 7 describes in detail the existing benchmark datasets including both classification and regression tasks. Finally, Section 8 sheds some light on the future research directions of SSL4NS-TD, and Section 9 concludes this survey.
Category Algorithm Encoder P-F Dataset Downstream Venue Code Link (Github)
Predictive Learning VIME${}^{[1]}$ MLP Both NeurIPS-20 jsyoon0823/VIME
TabTransformer${}^{[2]}$ Transformer Classification / lucidrains/tab-transformer-pytorch
TabNet${}^{[3]}$ Transformer Both AAAI-21 dreamquark-ai/tabnet
SEFS${}^{[4]}$ MLP Classification ICLR-22 chl8856/SEFS
LIFT${}^{[5]}$ Transformer Both NeurIPS-22 UW-Madison-Lee-Lab/LanguageInterfacedFineTuning
TapTap${}^{[6]}$ Transformer Both EMNLP-23 ZhangTP1996/TapTap
TabPFN${}^{[7]}$ Transformer Classification ICLR-23 automl/TabPFN
GReaT${}^{[8]}$ Transformer Both ICLR-23 kathrinse/be_great
STUNT${}^{[9]}$ MLP Classification ICLR-23 jaehyun513/STUNT
LFR${}^{[10]}$ MLP Classification NeurIPS-23 TRL layer6ai-labs/lfr
SwitchTab${}^{[11]}$ Transformer Classification AAAI-24 /
TP-BERTa${}^{[12]}$ Transformer Both ICLR-24 /
Contrastive Learning SCARF${}^{[13]}$ MLP Classification ICLR-22 clabrugere/pytorch-scarf
STab${}^{[14]}$ MLP Classification NeurIPS-22 TRL /
TransTab${}^{[15]}$ Transformer Classification NeurIPS-22 RyanWangZf/transtab
PTaRL${}^{[16]}$ MLP Both ICLR-24 /
Hybrid Learning SubTab${}^{[17]}$ MLP Classification NeurIPS-21 AstraZeneca/SubTab
SAINT${}^{[18]}$ Transformer Both NeurIPS-22 TRL somepago/saint
ReConTab${}^{[19]}$ Transformer Classification NeurIPS-22 TRL /
/${}^{[20]}$ Both Classification ICLR-23 LevinRoman/tabular-transfer-learning
DoRA${}^{[21]}$ MLP Regression CIKM-23 wwweiwei/DoRA
CT-BERT${}^{[22]}$ Transformer Classification / /
XTab${}^{[23]}$ Transformer Both ICML-23 BingzhaoZhu/XTab
UniTabE${}^{[24]}$ Transformer Both ICLR-24 /
${}^{[1]}$oon et al. (2020), ${}^{[2]}$uang et al. (2020), ${}^{[3]}$rik and Pfister (2021), ${}^{[4]}$ee et al. (2022), ${}^{[5]}$inh et al. (2022), ${}^{[6]}$hang et al. (2023), ${}^{[7]}$ollmann et al. (2023), ${}^{[8]}$orisov et al. (2023), ${}^{[9]}$am et al. (2023), ${}^{[10]}$ui et al. (2023), ${}^{[11]}$u et al. (2024), ${}^{[12]}$nonymous (2024a), ${}^{[13]}$ahri et al. (2022), ${}^{[14]}$ajiramezanali et al. (2022), ${}^{[15]}$ang and Sun (2022), ${}^{[16]}$nonymous (2024b), ${}^{[17]}$car et al. (2021), ${}^{[18]}$omepalli et al. (2022), ${}^{[19]}$hen et al. (2023), ${}^{[20]}$evin et al. (2023), ${}^{[21]}$u et al. (2023), ${}^{[22]}$e et al. (2023), ${}^{[23]}$hu et al. (2023), ${}^{[24]}$ang et al. (2023)
@@
2 Overview
2.1 Problem Definiton
Tabular data consist of horizontal rows as samples and vertical columns as features of the corresponding samples, where samples and features are structured in a tabular form. The features can be depicted with numbers, indicators, categories, text, etc. The applications of the tabular domain can be formulated into two categories: classification and regression tasks.
Formally, given a tabular dataset $D=\{x_{i},y_{i}\}_{i=1}^{|D|}$ with $|D|$ samples, where $x_{i}\in X\in\mathbb{R}^{N}$ denotes a sample consisting of $N$ features, the corresponding label $y_{i}\in Y$ is a scalar for regression tasks or is a class for classification tasks. The goal is to learn a predictive model $f:X\rightarrow Y$ trained by supervised loss function (e.g., cross-entropy or MSE). When applying SSL for tabular applications, the first intention is to construct an encoder function $e:X\rightarrow Z$, where $Z$ represents contextualized representations learned from self-supervised objectives. The encoder function can then be used with the model in the downstream task to utilize better representations to predict the label, which is often trained either in a two-stage (i.e., pre-train then fine-tune) manner or jointly with self-supervised and downstream objectives.
2.2 Taxonomy
Although relatively less concentrated work has been done in SSL4NS-TD compared with other domains, various SSL4NS-TD methods have been proposed to advance the exploration of implicit tabular structures without labels, which motivates us to outline the advancements of these works. To better understand the developing venation of SSL4NS-TD, we identify representative and profound research works in top-influential venues as well as high-impact but preprint or workshop papers, analyze their research motivations, and summarize their key technical contributions. Since there is no survey literature summarizing the efforts of SSL4NS-TD, this survey establishes a novel taxonomy of SSL4NS-TD that categorizes the existing research works into three major SSL groups as presented in Table 1. We briefly explain the core ideas of the three SSL research trends in NS-TD as follows.
@@ 2.2 Taxonomy
  • Predictive Learning of SSL4NS-TD is the most widely-used category for SSL4NS-TD to benefit the downstream performance. Due to the heterogeneous characteristics of features, designing prediction tasks before the final goal (i.e., downstream task) enables the model to learn background knowledge from raw data. The difficulty resides in devising predictive pretext tasks that are effective, considering the relations between upstream and downstream datasets and tasks. Although there is no consensus of designing predictive pretext tasks, several paradigms are proposed, including learning from masked features Huang et al. (2020); Yoon et al. (2020); Arik and Pfister (2021); Lee et al. (2022); Wu et al. (2024), perturbation in latent space Nam et al. (2023); Sui et al. (2023), and inherent in pre-trained language models Dinh et al. (2022); Borisov et al. (2023); Zhang et al. (2023); Anonymous (2024a).
@@
  • Contrastive Learning of SSL4NS-TD aims to learn the similarities and discrepancies of instances in the tabular domain. The main advantage is that contrastive learning offers a task-agnostic learning strategy that can be applied in a wide range of downstream applications and transferability with only a few labeled samples; however, the challenge lies in the design of which instances should be closer and be pulled over. Researchers thus address it by adopting different views of tabular data, such as instance-wise Bahri et al. (2022), model-wise Hajiramezanali et al. (2022), column-wise Wang and Sun (2022), and latent space-wise Anonymous (2024b).
@@
  • Hybrid Learning of SSL4NS-TD extends to combine predictive learning and contrastive learning as the SSL objective, which is moving towards a more unified SSL4NS-TD. The principal benefit of hybrid learning is that it integrates the advantages of both learning strategies. As there are various successful paradigms of predictive learning, researchers have attempted to explore the effectiveness of SSL4NS-TD by the combination of perturbation and contrastive learning Ucar et al. (2021); Somepalli et al. (2022); Chen et al. (2023); Yang et al. (2023); Zhu et al. (2023) as well as masking and contrastive learning Du et al. (2023); Levin et al. (2023); Ye et al. (2023).
@@ 3 Predictive Learning of SSL4NS-TD
As the compositions of NS-TD are heterogeneous yet without explicit relations (i.e., each column serves as a unique feature), it is challenging to distinguish relations from tables, which is one of the reasons that tree-based models are superior Grinsztajn et al. (2022). Motivated by SSL works in the homogeneous feature type domains (e.g., text, audio, image), such as perturbation, rotation, cropping, and adding noise, as predictive pretext tasks Balestriero et al. (2023), tabular-based SSL objectives are mainly designed on top of these approaches. The model is expected to be effective in downstream tasks if it is able to infer the original feature from the other masked or corrupted features. Formally, a general predictive model can be defined as:
@@
where $\psi$ refers to the loss function that optimizes transformed input $x^{*}_{i}$ with the self-supervised label $y^{*}_{i}$, which are created by the transformed function $\delta$. $g$ stands for the projection head aiming to convert encoded embeddings into the self-supervised prediction.

### v3

@@ Abstract
Self-supervised learning (SSL) has been incorporated into many state-of-the-art models in various domains, where SSL defines pretext tasks based on unlabeled datasets to learn contextualized and robust representations. Recently, SSL has become a new trend in exploring the representation learning capability in the realm of tabular data, which is more challenging due to not having explicit relations for learning descriptive representations. This survey aims to systematically review and summarize the recent progress and challenges of SSL for non-sequential tabular data (SSL4NS-TD). We first present a formal definition of NS-TD and clarify its correlation to related studies. Then, these approaches are categorized into three groups – predictive learning, contrastive learning, and hybrid learning, with their motivations and strengths of representative methods in each direction. Moreover, application issues of SSL4NS-TD are presented, including automatic data engineering, cross-table transferability, and domain knowledge integration. In addition, we elaborate on existing benchmarks and datasets for NS-TD applications to analyze the performance of existing tabular models. Finally, we discuss the challenges of SSL4NS-TD and provide potential directions for future research. We expect our work to be useful in terms of encouraging more research on lowering the barrier to entry SSL for the tabular domain, and of improving the foundations for implicit tabular data.
keywords:
Tabular Data, Self-Supervised Learning, Non-Sequential Tables
@@ 1 Introduction
Supervised learning has shown outstanding performance on various machine learning tasks; however, its main hurdles lie in heavily depending on expensive human annotations and generalization bottlenecks. With the scaling of the accessibility to unlabeled data, self-supervised learning (SSL) has demonstrated its generic and robust ability of learning contextualized information from correlations within the data. A line of studies has shown that SSL is able to push new boundaries in various domains, including text [18, 26], vision [27, 42], and speech [3, 44]. In addition, SSL illustrates strong generalizability, enabling models to adapt to tasks with limited labeled records and even unseen tasks [5]. The key advantage of SSL is that it reduces efforts in annotating a large amount of data, while further providing generalization ability.
Figure 1: Overall pipeline of SSL4NS-TD. Given tabular data, the SSL4NS-TD approaches adopt predictive learning (§$\S$3), contrastive learning (§$\S$4), or hybrid learning (§$\S$5) as the self-supervised objective before supervised fine-tuning on the downstream applications. Then, the trained model is evaluated based on the demand-related benchmarks (§$\S$7), which are framed as classification or regression problems.
@@
Previous SSL methods highly relied on the unique structure of the domain datasets, such as spatial relationships in images, semantic relationships in text, and vocal relationships in speech. In contrast, tabular data have no explicit relation between each feature, and may be completely different across various tabular datasets. Figure 1 describes the overall pipeline of learning representations from tabular data. To opt for the learning strategy, in contrast to only using supervised learning requiring task-specific labels during the training stage, SSL is further leveraged to learn task-agnostic representations from pretext tasks, creating labels explicitly (i.e., predictive learning) and/or implicitly (i.e., contrastive learning) as categorized by [28]. The model is expected to learn universal representations from unlabeled tabular datasets and accordingly adapt effectively to different downstream tasks including both classification and regression problems. Generally, existing techniques with SSL for tabular data can be grouped into sequential and non-sequential tabular types for representation learning.
@@
This survey focuses on SSL for non-sequential tabular data (SSL4NS-TD) for two reasons. First, SSL for sequential data commonly used in fields such as recommendation systems has been widely adopted by recurrent- and Transformer-based techniques based on the temporal-ordered composition, and has already been discussed in recent survey papers [2, 17]. On the other hand, non-sequential tabular data (NS-TD) make it more challenging to define explicit structures with pretext tasks, which has led to tremendous efforts in recent years, but has not yet been discussed in the literature. This inspired us to provide a systematic review of recent SSL4NS-TD approaches to discuss their motivations and self-supervised objectives. Second, whether deep learning models are superior to machine learning models in the NS-TD problems remains an active discussion topic [15, 29]. However, SSL showcases its robustness in not only full-labeled data but also only a few labeled records, where machine learning and deep learning models fall short in such low-resource scenarios. It is thus urgent to summarize the efforts of learning contexts with self-supervision for NS-TD to apprise the community of the current progress and to open more discussion.
The main contributions of our survey are as follows:
  • We present the first comprehensive survey of recent advancements in SSL4NS-TD, consisting of problem definitions, taxonomy, application issues, NS-TD datasets, and evaluation protocols. Our systematic categorization of these methods into three subcategories based on their defining characteristics offers an in-depth and structured overview of the subject. Moreover, this survey presents an up-to-date paper survey with a paper list that will be continuously updated [^1] A detailed list can be found at https://github.com/wwweiwei/awesome-self-supervised-learning-for-tabular-data..
  • We empirically evaluate representative SSL4NS-TD approaches of each learning category on the most recent and large-scale benchmark, TabZilla [29], validating the most state-of-the-art SSL4NS-TD methods. In addition, we delve into these comparisons to provide insights into stimulating future advancements and the gap across various learning categories.
  • We highlight the addressed challenges and future directions in SSL for the existing NS-TD methods. We intend to shed light on under-researched aspects to spur further investigations that will explore the various possible paths and their interrelations within NS-TD applications, e.g., finance and healthcare.
@@
The remainder of this review is structured as follows. Section 2 first introduces the problem of NS-TD and proposes a novel taxonomy of SSL4NS-TD consisting of three learning strategies as summarized in Table 1. The detailed discussion of the achievements, the downstream tasks, and code links (if available) of the three directions are summarized in Sections 3, 4, and 5 respectively. Section 6 briefly introduces application issues (i.e., automatic data engineering, cross-table transferability, and domain knowledge integration) in practice and recent related work. Subsequently, Section 7 describes in detail the existing benchmark datasets including both classification and regression tasks, and the benchmark results on several SSL4NS-TD approaches are reported and analyzed. Finally, Section 8 sheds some light on the future research directions of SSL4NS-TD, and Section 9 concludes this survey.
2 Overview
2.1 Problem Definition of SSL4NS-TD
Tabular data consist of horizontal rows as samples and vertical columns as features of the corresponding samples, where samples and features are structured in a tabular form. The features can be depicted with numbers, indicators, categories, text, etc. The applications of the tabular domain can be generally formulated into two categories: classification and regression tasks.
In the definition of the NS-TD tasks, given a tabular dataset $D=\{x_{i},y_{i}\}_{i=1}^{|D|}$ with $|D|$ samples, where $x_{i}\in X\in\mathbb{R}^{N}$ denotes a sample consisting of $N$ features, the corresponding label $y_{i}\in Y$ is a scalar for regression tasks or is a class for classification tasks. The goal is to learn a predictive model $f:X\rightarrow Y$ trained by supervised loss function (e.g., cross-entropy or MSE). Distinct from sequential tabular data, non-sequential tabular data do not have explicit chronological order (e.g., timestamp) between samples. When applying SSL for NS-TD applications, the first intention is to construct an encoder function $e:X\rightarrow Z$, where $Z$ represents contextualized representations learned from self-supervised objectives. It is worth noting that the labels from self-supervised objectives are obtained from the data itself instead of requiring manual annotations [28]. The encoder function can then be used with the model in the downstream task to utilize better representations to predict the label, which is often trained either in a two-stage (i.e., pre-train then fine-tune) manner or jointly with self-supervised and downstream objectives.
2.2 Taxonomy
Although relatively less concentrated work has been done in SSL4NS-TD compared with other domains, various SSL4NS-TD methods have been proposed to advance the exploration of implicit tabular structures without labels, which motivated us to outline the advancements of these works. To better understand the developing venation of SSL4NS-TD, we have identified representative and profound research works in the most influential venues as well as high-impact but preprint or workshop papers, analyzed their research motivations, and summarized their key technical contributions. Since there is no survey literature summarizing the efforts of SSL4NS-TD, this survey establishes a novel taxonomy of SSL4NS-TD that categorizes the existing research works into three major SSL groups, as presented in Table 1. We briefly explain the core ideas of the three SSL research trends in NS-TD as follows.
Category Algorithm Encoder P-F Dataset Downstream Venue Code Link (Github) Predictive Learning VIME[1] MLP Both NeurIPS-20 jsyoon0823/VIME TabTransformer[2] Transformer Classification / lucidrains/tab-transformer-pytorch TabNet[3] Transformer Both AAAI-21 dreamquark-ai/tabnet SEFS[4] MLP Classification ICLR-22 chl8856/SEFS LIFT[5] Transformer Both NeurIPS-22 UW-Madison-Lee-Lab/LanguageInterfacedFineTuning TapTap[6] Transformer Both EMNLP-23 ZhangTP1996/TapTap TabPFN[7] Transformer Classification ICLR-23 automl/TabPFN GReaT[8] Transformer Both ICLR-23 kathrinse/be_great STUNT[9] MLP Classification ICLR-23 jaehyun513/STUNT SwitchTab[10] Transformer Classification AAAI-24 / LFR[11] MLP Classification ICLR-24 layer6ai-labs/lfr TP-BERTa[12] Transformer Both ICLR-24 / Contrastive Learning SCARF[13] MLP Classification ICLR-22 clabrugere/pytorch-scarf STab[14] MLP Classification NeurIPS-22 TRL / TransTab[15] Transformer Classification NeurIPS-22 RyanWangZf/transtab PTaRL[16] MLP Both ICLR-24 / Hybrid Learning SubTab[17] MLP Classification NeurIPS-21 AstraZeneca/SubTab SAINT[18] Transformer Both NeurIPS-22 TRL somepago/saint ReConTab[19] Transformer Classification NeurIPS-22 TRL / /[20] Both Classification ICLR-23 LevinRoman/tabular-transfer-learning DoRA[21] MLP Regression CIKM-23 wwweiwei/DoRA CT-BERT[22] Transformer Classification / / XTab[23] Transformer Both ICML-23 BingzhaoZhu/XTab UniTabE[24] Transformer Both ICLR-24 / [1][50], [2][20], [3][1], [4][23], [5][12], [6][51], [7][19], [8][8], [9][30], [10][43], [11][36], [12][46], [13][4], [14][16], [15][40], [16][49], [17][38], [18][35], [19][9], [20][25], [21][13], [22][48], [23][53], [24][47]
@@ 2.2 Taxonomy
  • Predictive Learning of SSL4NS-TD is the most widely-used category for SSL4NS-TD to benefit the downstream performance. Due to the heterogeneous characteristics of features, designing prediction tasks before the final goal (i.e., downstream task) enables the model to learn background knowledge from raw data. The difficulty resides in devising predictive pretext tasks that are effective, considering the relations between upstream and downstream datasets and tasks. Although there is no consensus of designing predictive pretext tasks, several paradigms have been proposed, including learning from masked features [20, 50, 1, 23, 43], perturbation in latent space [30, 36], and inherent in pre-trained language models [12, 8, 51, 46].
@@
  • Contrastive Learning of SSL4NS-TD aims to learn the similarities and discrepancies of instances in the tabular domain. The main advantage is that contrastive learning offers a task-agnostic learning strategy that can be applied in a wide range of downstream applications and transferability with only a few labeled samples; however, the challenge lies in the design of which instances should be closer and be pulled over. Researchers thus address it by adopting different views of tabular data, such as instance-wise [4], model-wise [16], column-wise [40], and latent space-wise [49].
@@
  • Hybrid Learning of SSL4NS-TD extends to combine predictive learning and contrastive learning as the SSL objective, which is moving towards a more unified SSL4NS-TD. The principal benefit of hybrid learning is that it integrates the advantages of both learning strategies. As there are various successful paradigms of predictive learning, researchers have attempted to explore the effectiveness of SSL4NS-TD by the combination of perturbation and contrastive learning [38, 35, 9, 47, 53] as well as masking and contrastive learning [13, 25, 48].
@@ 3 Predictive Learning of SSL4NS-TD
As the compositions of NS-TD are heterogeneous yet without explicit relations (i.e., each column serves as a unique feature), it is challenging to distinguish relations from tables, which is one of the reasons that tree-based models are superior [15]. Motivated by SSL works in the homogeneous feature type domains (e.g., text, audio, image), such as perturbation, rotation, cropping, and adding noise, as predictive pretext tasks [5], tabular-based SSL objectives are mainly designed on top of these approaches to design pretext tasks. The model is expected to be effective in downstream tasks if it is able to infer the original feature from the other masked or corrupted features. Formally, a general SSL predictive model can be defined as:
@@
where $\psi$ refers to the loss function that optimizes transformed input $x^{*}_{i}$ with the self-supervised label $y^{*}_{i}$, which are created by the transformed function $\delta$. $g$ stands for the projection head aiming to convert encoded embeddings by encoder $e$ into the self-supervised prediction.